# Geometric integrators for adiabatically closed simple thermodynamic systems

### v2

@@ 4.4 Expansion of a Van der Waals gas against external pressure
To make a comparison of the results given by the discrete thermodynamic forced Euler-Lagrange equations and the ones obtained via the GENERIC-Motivated method introduced in [Fulop2025], we will work with the reduced version of the Van der Waals equations. The continuous Lagrangian for this system, considering a monoatomic gas with three degrees of freedom, is given by:
@@
where $m_{\mathrm{gas}}$ is the mass of a gas molecule. We have assumed that the gas in the piston is He-4 and used the values of its critical magnitudes and molar mass in [He41989]. From now on, we will consider, as in [Fulop2025], the quotient $\frac{m_{p}}{A^{2}}=1$.
@@
The results have been compared with the ones given by the GENERIC-motivated extended symplectic method considered in [Fulop2025] (choosing the coefficients $\alpha_{b}=\alpha_{p}=1$ for such method), which was shown to outperform the classical Euler method.
@@ 5 Conclusions and outlook
Inspired by the geometric framework of almost cosymplectic manifolds derived in [d.B2025], we have developed a discrete variational principle for adiabatically closed simple thermodynamic systems, leading to the discrete thermodynamic forced Euler–Lagrange equations (22). We have also presented a Noether’s theorem for these systems (Theorem 13) as well as for the continuous counterpart (Theorem 3 and Theorem 5. We have illustrated the effectiveness of our method with four examples: the damped harmonic oscillator, both a perfect gas and a Van der Waals gas confined by a piston contained in a cylinder, and a Van der Waals gas subject to external forces.
@@
  • To extend these methods to the remaining thermodynamic systems considered in [G.Y2018, Y.G2022], namely systems with internal mass transfer, adiabatically closed non-simple thermodynamic systems, and open simple thermodynamic systems.
  • To investigate the relationship between thermodynamic systems and nonholonomic constraints (see [lucia]). Furthermore, we would like to relate our approach to the symplectic one in [ghosh2024hamiltonianthermodynamicssymplecticmanifolds].

### v3

@@ 4.4 Expansion of a Van der Waals gas against external pressure
To make a comparison of the results given by the discrete thermodynamic forced Euler-Lagrange equations and the ones obtained via the GENERIC-Motivated method introduced in [14], we will work with the reduced version of the Van der Waals equations. The continuous Lagrangian for this system, considering a monoatomic gas with three degrees of freedom, is given by:
@@
where $m_{\mathrm{gas}}$ is the mass of a gas molecule. We have assumed that the gas in the piston is He-4 and used the values of its critical magnitudes and molar mass in [4]. From now on, we will consider, as in [14], the quotient $\frac{m_{p}}{A^{2}}=1$.
@@
The results have been compared with the ones given by the GENERIC-motivated extended symplectic method considered in [14] (choosing the coefficients $\alpha_{b}=\alpha_{p}=1$ for such method), which was shown to outperform the classical Euler method.
@@ 5 Conclusions and outlook
Inspired by the geometric framework of almost cosymplectic manifolds derived in [10], we have developed a discrete variational principle for adiabatically closed simple thermodynamic systems, leading to the discrete thermodynamic forced Euler–Lagrange equations (22). We have also presented a Noether’s theorem for these systems (Theorem 13) as well as for the continuous counterpart (Theorem 3 and Theorem 5. We have illustrated the effectiveness of our method with four examples: the damped harmonic oscillator, both a perfect gas and a Van der Waals gas confined by a piston contained in a cylinder, and a Van der Waals gas subject to external forces.
@@
  • To extend these methods to the remaining thermodynamic systems considered in [16, 33], namely systems with internal mass transfer, adiabatically closed non-simple thermodynamic systems, and open simple thermodynamic systems.
  • To investigate the relationship between thermodynamic systems and nonholonomic constraints (see [26]). Furthermore, we would like to relate our approach to the symplectic one in [18].